Problem #1. The difference between 2 numbers are 8. Twice the smaller plus 3 times the larger is 109.
Problem #2. The perimeter of a standard basketball court is 288 ft. Find the dimensions of the basketball court if the length is 4 ft. longer than the width?
Please help and would appreciate if you can solve it with complete solutions and checking for me to understand it better...I'm having problems solving this kind step by step...thank you.

Answer by htmentor(1343)

Problem #1. The difference between 2 numbers are 8. Twice the smaller plus 3 times the larger is 109.
===============================
Let x = one of the numbers
Then the other number is x - 8
2(x-8) + 3x = 109
Solve for x:
2x - 16 + 3x = 109
5x = 125
x = 25
So the two numbers are 25, 17
Problem #2. The perimeter of a standard basketball court is 288 ft. Find the dimensions of the basketball court if the length is 4 ft. longer than the width?
===============================
Perimeter of a rectangle = 2w + 2l, where l is the length, w is the width
Given: l = w + 4, perimeter = 288
Substitute the value for l in the equation for the perimeter:
2w + 2(w+4) = 288
4w + 8 = 288
w = 280/4 = 70
Ans: width = 70 ft, length = 74 ft.